

********************************************************************
UARS WEEKLY STATUS REPORT - 12 JUL 2002
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, taking solar data on all
available orbital tracking periods.
***********
SUSIM
***********
Noting received.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
PEM continues to work with the GUVI team on the diagnosis of data from
the first conjunction period. Analysis of the two conjunction passes showed
that the aurora as detected by GUVI was at higher latitude than UARS.
PEM has been working with the GUVI team to identify a second time period
for coordinated measurements. The orbital conditions are different from the
first period where now the TIMED-UARS orbits are nearly perpendicular.
PEM operation is requested to begin three orbits prior to and last for three
orbits after conjunction times. Predicted orbital conjunctions as of July 5
(varies with each new set of predictions) are:
STK
Contact # Date Day Time Geo Lat ILAT MLT Hemisphere Note
(deg) (deg) (hr)
9 July 8 189 06:21 -53.33 55.45 7.17 South
10 July 8 189 07:09 53.23 51.23 19.63 North
11 July 8 189 07:57 -53.12 52.70 7.33 South
12 July 13 194 00:34 51.54 62.04 19.92 North
13 July 13 194 01:22 -51.41 63.44 6:37 South
14 July 13 194 02:11 49.34 59.59 19.55 North
15 July 17 198 19:36 48.20 45.47 19.17 North
16 July 17 198 20:24 -48.05 59.76 7.26 South
17 July 17 198 21:12 47.91 50.81 19.49 North
As dates of predicted orbital elements improve, contacts appear and disappear
from this list; however, in general the conjunctions seem to be occurring at
higher latitudes later in the measurement time period (higher latitude
contacts increase the chance for adequate precipitation to produce a common
signal), This is opposite to the early contact predictions for this conjunction
time period.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered. NOAA has changed the binary structure of the NMC data and
PEM is now receiving two data streams, one in the original VAX format and the
other from their new binary format.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG attended the UARS Monthly Status Review Meeting
on July 10, 2002.
The MPG continues to serve as the point-of-contact for UARS
instrument operations to support re-engineering of the UARS
Mission Operations Concept to develop capabilities for the
Science Traceability Mission.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 29 June 2002 (Orbit 59046, GMT Day 180) through
05 July 2002 (Orbit 59151, GMT Day 186).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
SUSIM 182/2149 59090 Power Off
SOLSTICE 182/2149 59090 Day-Only
PEM 182/2149 59090 Power On
HALOE 186/1459 59142 Power Off
The beta angle ranged between the angles of 26.0 degrees to 55.3
degrees for this report period. The beta angle is now increasing
towards a maximum angle of 79 degrees on 13 July 2002 (DOY 194).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.8 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 0.63 to 1.25 degrees C. Additional
battery performance data for each battery on 21 June and 28 June 2002
are provided in Appendix D.
The clock error ranged from - 10.7 to 10.2 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 01
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 02
HALOE = 02
SUSIM = 01
SOLSTICE = 01
PEM = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 02
One (1) Anomaly Report remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Two (2) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT was able to get IP NOCC/WSGT to change the firewall rules for
the new TAC systems so that TACs 4 & 5 could be moved to the control
center in Bldg 32. Test commands were successfully sent from TAC 3
and TAC 4 after the reconfiguration to insure their proper execution.
The ground system is now flying on TAC 4 for operations with TAC 5 as
a backup.
MEETINGS
-None this report period
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 02 August 2002 (DOY 214).
This maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 05 July 2002
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
59089 182/2052 02-227 HALOE Scan Commands
59090 182/2149 02-228 SUSIM Power Off
59090 182/2149 02-229 SOLSTICE Day-Only Ops
59090 182/2149 02-230 PEM ZEPS/HEPS Fulltime
59121 185/0008 02-231 OBC Clk Adj from 0.00 to 20.6 msec/day
59142 186/0942 02-232 HALOE Power Off (TBAT)
59146 186/1459 02-233 OBC Clk Adj from 20.6 to 0.00 msec/day
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 05 July 2002
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
HRPTSBARTEMP Y-HI HRDI Barrel Temp 58250-64
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
______________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 05 July 2002
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
59121 185/0008 CLK ADJ FR 0.00 TO 20.6 MSEC/DAY
59146 186/1459 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 05 July 2002
28 June 2002 - GMT Day 179 Beta = 22.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 +728.0/+728.0 -3.68/-4.31 +00.0/+00.0 9.4 0.000 0.0
2 +05.6/-078.4 +5.15/+4.51 +27.2/-14.4 25.1 1.030 20.2
3 +16.8/-022.4 +4.19/+3.23 +26.8/-14.4 25.1 1.020 20.0
05 July 2002 - GMT Day 186 Beta = 55.3 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 +33.6/-22.4 -3.05/-4.62 +00.0/+00.0 8.5 0.000 0.0
2 00.0/-69.6 +4.50/+2.60 +22.0/-16.4 25.8 1.040 17.8
3 +22.4/-22.4 +3.55/+1.66 +21.7/-16.2 25.8 1.035 17.7
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period
APPENDIX E
UARS ANOMALY REPORTS
Week ending 05 July 2002
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-039 FOT See List July 30832 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-033 FOT 58553 147/0359 N/A SSPP IN WAIT MODE/PLANNING ERR
02-035 FOT 58719 158/0606 44490 WSGT EQUIP PROBLEMS
02-035 FOT 58720 158/0746 44490 WSGT EQUIP PROBLEMS
02-036 FOT See List June 01 N/A Generic Late Acquisition
02-037 FOT 58927 172/0223 N/A Data Dropout CDS#30824
02-038 FOT 58959 174/0526 N/A LATE ACQ EQUIP PROB at WS
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
02-039 FOT See List July 30832 Generic Late Acquisition